# JAFAR: Jack up Any Feature at Any Resolution

### v1

@@ 4.2 Qualitative Comparisons
To qualitatively evaluate the upsampled feature maps produced by various baselines, we project all features onto a shared 3-dimensional PCA basis, mapping them into a common RGB space. As shown in Figs. 3 and 5, the low-resolution features—due to the spatial compression imposed by the vision encoder’s patch size—reveal large, blocky regions that capture semantic content but fail to preserve fine image geometry, object boundaries, or shape details. Bilinear upsampling, which interpolates features without considering image content, yields blurry output feature maps that preserve positional embeddings artifacts without adding meaningful detail. While methods like Large-Image and Strided preserve sharpness, their outputs are noisier and less coherent than JAFAR’s. Furthermore, they are more computationally demanding, as they require the vision encoder to process a larger number of patches (see Tab. 10). JAFAR shows a clear qualitative advantage over all baselines, consistently producing sharp features that accurately capture image structure. It is also the only task-agnostic method that effectively suppresses artifacts from positional embeddings in the low-resolution features.
@@ 4.3 Transfer on Downstream Tasks
Since upsampled features are expected to provide a richer signal for downstream tasks, we evaluate their effectiveness on two benchmarks: linear-probing semantic segmentation and depth estimation, using DINOv2 ViT-S/14 as the foundation vision encoder. For the Large-Image and Strided baselines, upsampling is performed during the encoder’s forward pass and followed by bilinear interpolation to reach the target output resolution. For task-agnostic upsamplers such as LiFT, FeatUp, and JAFAR, we pre-train the upsampling module on the corresponding backbone, then freeze it and apply it after feature extraction. The linear probe is trained independently of the upsampler. For task-dependent methods—including CARAFE, SAPA, ReSFu, and DySample—we jointly train both the upsampler and the linear probe on each dataset and task. All experiments (except Large-Image) use input images of resolution $448\times 448$, with target labels at the same resolution.
@@ 4.3.1 Semantic Segmentation
For semantic segmentation, we train a linear projection head to predict coarse class labels using a cross-entropy loss across several benchmark datasets: COCO-Stuff [40] (27 classes), ADE20K [41] (150 classes), Pascal VOC [42] (21 classes including background), and Cityscapes [43] (27 classes). The linear layer is trained for 5 epochs on COCO-Stuff and 20 epochs on the remaining datasets, using a batch size of 4. Performance is evaluated on the respective validation sets using mean Intersection-over-Union (mIoU) and pixel-wise accuracy.
@@
DINOv2-ViT-S/14 Semantic Segmentation Depth Estimation COCO VOC ADE20K Cityscapes COCO mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) $\delta_{1}$ ($\uparrow$) RMSE ($\downarrow$) Training-free    Nearest 56.17 76.97 76.41 93.80 37.27 71.91 54.05 90.36 58.08 0.70    Bilinear 59.03 79.07 80.70 95.17 39.23 73.69 59.37 92.47 59.92 0.66    Large Image (x8) 56.94 88.60 26.42 66.39 47.72 92.49    Strided 55.93 77.40 75.88 93.94 36.15 72.08 59.26 92.57 56.98 0.70 Task-Dependent    CARAFE [22] 59.73 79.65 80.26 95.14 38.30 73.42 56.05 91.83 61.42 0.64    SAPA [23] 57.77 78.28 77.02 94.07 35.87 71.85 50.12 90.02 60.34 0.67    DySample [24] 59.50 79.42 81.62 95.48 38.99 73.62 59.71 92.69 61.25 0.64    ReSFU [25] 60.08 79.84 80.30 95.05 38.91 73.93 55.53 91.62 66.14 0.56 Task-Agnostic    FeatUp [28] 60.10 79.95 81.08 95.32 38.82 73.74 56.06 91.86 61.69 0.64    LIFT [27] 58.18 78.95 78.06 94.62 38.73 73.69 58.75 92.60 57.04 0.70    JAFAR 60.78 80.47 84.44 96.28 40.49 74.92 61.47 93.42 62.18 0.62
As shown in Tab. 1, JAFAR consistently achieves the highest performance across all four semantic segmentation benchmarks, in both mIoU and accuracy. On average, JAFAR delivers a $+1.63$ mIoU improvement over the next-best method across all datasets. Compared to FeatUp, JAFAR achieves an average gain of $+2.78$ mIoU corresponding to a $+4.8\%$ gain, with a peak improvement of $+5.41$ mIoU ($+9.7\%$) on Cityscapes. Fig. 4 shows linear probe segmentation result.
@@ 4.3.2 Depth Estimation
For depth estimation, we follow the approach in [28] and train on pseudo-labels generated by the state-of-the-art Depth Anything V2 network [16]. We report two standard metrics from the monocular depth estimation literature: root mean square error (RMSE) and $\delta_{1}<1.25$. The $\delta_{1}$ metric measures the percentage of pixels where the predicted depth $y$ is within 25% of the ground-truth $y^{*}$, formally defined as $\delta_{1}=max\left(\frac{y}{y^{*}},\frac{y^{*}}{y}\right)<1.25$. We train the linear probe for 5 epochs on the COCO training set, using a batch size of 4. Although JAFAR was not trained on this specific task, we observe that it reaches competitive scores, ranking second among the baselines. Notably, JAFAR outperforms both FeatUp and LiFT while also surpassing all task-dependent methods but ReSFU. Fig. 4 shows linear probe depth estimation result.
@@ 4.3.3 Class Activation Maps Faithfulness
Following the approach in [28], our method can be seamlessly integrated into explainability tools such as Class Activation Maps (CAMs). Despite recent advances, CAMs are still fundamentally limited by the low-resolution feature maps produced by standard vision encoders, which hinders their ability to localize fine-grained details. By upsampling the features, our method yields sharper and more informative explanations. To assess the quality of the resulting CAMs, we adopt standard evaluation metrics from the literature: Average Drop (A.D), Average Increase (A.I), Average Gain (A.G), Coherency (Coh.), and Complexity (Cplx.).
@@
A.D ($\downarrow$) A.I ($\uparrow$) A.G ($\uparrow$) Coh. ($\uparrow$) Cplx. ($\downarrow$) ADCC ($\uparrow$) Training-free    Bilinear 19.0 18.5 3.4 88.8 59.9 61.7 (-11.6)    Large Image (x8) 48.8 12.8 2.5 67.9 38.0 59.5 (-13.8)    Strided 19.8 15.1 3.5 85.9 54.2 65.3 (-8.0) Task-Dependent    CARAFE [22] 49.9 4.8 1.0 66.9 34.9 59.7 (-13.6)    SAPA [23] 8.5 32.7 4.1 96.5 69.6 55.4 (-17.9)    Dysample [24] 17.8 20.0 3.8 90.2 60.7 61.6 (-11.7)    ReSFU [25] 14.5 24.0 3.7 92.1 64.2 59.4 (-13.9) Task-Agnostic    FeatUp [28] 15.3 24.0 4.3 91.6 58.2 64.3 (-9.0)    LiFT [27] 66.9 8.7 2.3 65.2 9.4 53.0 (-20.3)    JAFAR 17.4 30.9 6.5 91.4 44.1 73.3
In particular, A.D, A.I, and A.G measure how sensitive the classifier’s output is to the most salient regions of the input—an effective CAM should highlight areas that, when masked, lead to a notable change in classification confidence. Since each of these metrics captures only a single aspect of CAM quality, we also report the ADCC score—an aggregate metric proposed in [44] that provides a more holistic evaluation. Additional details are provided in Supp. B. As illustrated qualitatively in Fig. 4, JAFAR generates sharper and more semantically accurate CAMs compared to all baselines. While training-free methods don’t help to recover important regions, task-dependent approaches typically produce blurrier and less precise maps. Quantitative results in Tab. 2 further support this, with JAFAR achieving the highest score on the aggregate ADCC metric—outperforming the second-best method by 8 points, a relative improvement of 12.5%.
@@ 4.3.4 Zero-Shot Open-Vocabulary Segmentation
We further evaluate our method on a zero-shot open-vocabulary segmentation task, following the setup from [9], where class labels from the dataset serve as textual inputs and predictions are made by selecting the class with the highest similarity score (argmax). Using a CLIP-ViT-B/16 backbone, this approach is entirely training-free, as it does not require a learned probing head. Results show that JAFAR significantly outperforms all baselines, with particularly strong improvements on Pascal VOC. Despite the increased difficulty of ADE20K, which includes 150 classes, our method still achieves the highest performance in both mIoU and accuracy. We report only FeatUp among the task-agnostic baselines, as it is the second-best performing method.
Table 3: Zero-Shot Open-Vocabulary Evaluation. Using MaskCLIP [9] for zero-shot open-vocabulary segmentation, JAFAR consistently improves performance, indicating strong alignment with the original features.
Upsampling VOC ADE20K Cityscapes mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) Nearest 24.13 30.80 9.33 24.65 19.66 50.27 Bilinear 27.87 35.27 11.03 27.78 21.56 53.21 Large Image (x2) 23.24 32.16 8.08 24.94 21.91 52.22 FeatUp [28] 32.27 39.78 13.03 33.28 24.76 60.11 JAFAR 35.70 44.93 13.61 33.28 25.26 61.73
@@ 4.3.5 Bird’s-Eye View Segmentation
Upsampling SimpleBeV [45] PointBeV [46] BeVFormer [47] Low-Res 31.75 34.89 33.72 Bilinear 33.67 36.01 34.18 FeatUp 33.95 35.38 34.01 JAFAR 36.59 37.20 36.54
Finally, we studied the impact of our upsampler in a complex training pipeline. The task takes several images taken from cameras as input and consists on outputting the bird’s-eye view (BeV) segmentation map. In our setup, we used a frozen DINOv2 [7] backbone and trained the rest of the architecture—namely, the upsampler, the BeV encoder, and the segmentation head. This task is particularly challenging, as the model must learn to map features from the image plane to the BeV plane. To ensure a fair comparison, we also trained the architecture without an upsampler, using lower-resolution input images (496×224). We adopted the optimization hyperparameters from PointBeV [46], adjusting the batch size to 1 and training for 100 epochs. Our results show that using an upsampler consistently improves predictions, regardless of the architecture employed—SimpleBev [45], PointBeV [46], or BevFormer [47]. Notably, performance improves significantly when using JAFAR as the upsampler, with mIoU gains up to $+5$ points.
@@ 4.4 Ablations
To evaluate the benefit of deriving both queries and keys from a shared image encoding, we compare in Tab. 5 several key generation strategies. In the Linear Projection baseline, keys are obtained by applying a simple linear layer to the low-resolution features $F_{lr}$ from the vision encoder, without using the image encoding. In the Concatenation baseline, we replace the modulation block with a direct concatenation of $F_{lr}$ and the preliminary keys $\tilde{K}$.
@@
Ablation Type / Setting Semantic Segmentation VOC ADE20K Cityscapes mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) Keys Strategy    Linear Projection 80.02 (-4.42) 94.87 (-1.41) 37.87 (-2.62) 73.22 (-1.70) 52.45 (-9.02) 90.80 (-2.62)    Concatenation 83.13 (-1.27) 95.94 (-0.34) 40.06 (-0.43) 74.56 (-0.36) 58.70 (-2.77) 92.65 (-0.77)    SFT Modulation 84.44 96.28 40.49 74.92 61.47 93.42 Attention Heads    $n=1$ 84.13 (-0.31) 96.21 (-0.07) 40.15 (-0.34) 74.79 (-0.13) 60.94 (-0.53) 93.32 (-0.10)    $n=2$ 84.27 (-0.17) 96.27 (-0.01) 40.42 (-0.07) 74.95 (+0.03) 61.19 (-0.28) 93.42 (-0.00)    $n=4$ 84.44 96.28 40.49 74.92 61.47 93.42    $n=8$ 83.82 (-0.62) 96.13 (-0.15) 40.07 (-0.42) 74.20 (-0.72) 60.56 (-0.91) 93.33 (-0.09)
Injecting semantic information into the keys through feature modulation is crucial for generating high-quality features and achieving strong query-key alignment. In comparison, the linear projection baseline shows a significant performance drop, and SFT consistently outperforms the concatenation approach. Increasing the number of attention heads up to 4 further enhances performance by producing more robust upsampling kernels through averaged post-softmax scores. Beyond this point, however, the benefits reverse: the per-head dimensionality becomes too low to support effective alignment, while the computational cost increases, ultimately degrading output quality.

### v2

@@ 4.2 Qualitative Comparisons
To qualitatively evaluate the upsampled feature maps produced by various baselines, we project all features onto a shared 3-dimensional PCA basis, mapping them into a common RGB space. As shown in Figs.˜3 and 5, the low-resolution features—due to the spatial compression imposed by the vision encoder’s patch size—reveal large, blocky regions that capture semantic content but fail to preserve fine image geometry, object boundaries, or shape details. Bilinear upsampling, which interpolates features without considering image content, yields blurry output feature maps that preserve positional embeddings artifacts without adding meaningful detail. While methods like Large-Image and Strided preserve sharpness, their outputs are noisier and less coherent than JAFAR’s. Furthermore, they are more computationally demanding, as they require the vision encoder to process a larger number of patches (see Tab.˜12). JAFAR shows a clear qualitative advantage over all baselines, consistently producing sharp features that accurately capture image structure. It is also the only task-agnostic method that effectively suppresses artifacts from positional embeddings in the low-resolution features.
@@ 4.3 Transfer on Downstream Tasks
Since upsampled features are expected to provide a richer signal for downstream tasks, we evaluate their effectiveness on two benchmarks: linear-probing semantic segmentation and depth estimation, using DINOv2 ViT-S/14 as the foundation vision encoder. For the Large-Image and Strided baselines, upsampling is performed during the encoder’s forward pass and followed by bilinear interpolation to reach the target output resolution. For task-agnostic upsamplers such as LiFT, FeatUp, and JAFAR, we pre-train the upsampling module on the corresponding backbone, then freeze it and apply it after feature extraction. The linear probe is trained independently of the upsampler. For task-dependent methods, including CARAFE, SAPA, ReSFu, and DySample, we jointly train both the upsampler and the linear probe on each dataset and task. All experiments (except Large-Image) use input images of resolution $448\times 448$, with target labels at the same resolution.
@@ 4.3.1 Semantic Segmentation
For semantic segmentation, we train a linear projection head to predict coarse class labels using a cross-entropy loss across several benchmark datasets: COCO-Stuff [lin2014microsoft] (27 classes), ADE20K [ade20k] (150 classes), Pascal VOC [pascalvoc] (21 classes including background), and Cityscapes [cityscapes] (27 classes). The linear layer is trained for 5 epochs on COCO-Stuff and 20 epochs on the remaining datasets, using a batch size of 4. Performance is evaluated on the respective validation sets using mean Intersection-over-Union (mIoU) and pixel-wise accuracy.
@@
DINOv2-ViT-S/14 Semantic Segmentation Depth Estimation COCO VOC ADE20K Cityscapes COCO mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) $\delta_{1}$ ($\uparrow$) RMSE ($\downarrow$) Training-free    Nearest 56.17 76.97 76.41 93.80 37.27 71.91 54.05 90.36 58.08 0.70    Bilinear 59.03 79.07 80.70 95.17 39.23 73.69 59.37 92.47 59.92 0.66    Large Image (x8) 56.94 88.60 26.42 66.39 47.72 92.49    Strided 55.93 77.40 75.88 93.94 36.15 72.08 59.26 92.57 56.98 0.70 Task-Dependent    CARAFE [carafe] 59.73 79.65 80.26 95.14 38.30 73.42 56.05 91.83 61.42 0.64    SAPA [sapa] 57.77 78.28 77.02 94.07 35.87 71.85 50.12 90.02 60.34 0.67    DySample [dysample] 59.50 79.42 81.62 95.48 38.99 73.62 59.71 92.69 61.25 0.64    ReSFU [resfu] 60.08 79.84 80.30 95.05 38.91 73.93 55.53 91.62 66.14 0.56 Task-Agnostic    FeatUp [featup] 60.10 79.95 81.08 95.32 38.82 73.74 56.06 91.86 61.69 0.64    LIFT [lift] 58.18 78.95 78.06 94.62 38.73 73.69 58.75 92.60 57.04 0.70    JAFAR 60.78 80.47 84.44 96.28 40.49 74.92 61.47 93.42 62.18 0.62
As shown in Tab.˜1, JAFAR consistently achieves the highest performance across all four semantic segmentation benchmarks, in both mIoU and accuracy. On average, JAFAR delivers a +1.63+1.63 mIoU improvement over the next-best method across all datasets. Compared to FeatUp, JAFAR achieves an average gain of +2.78+2.78 mIoU corresponding to a $4.8\%$ gain, with a peak improvement of +5.41+5.41 mIoU ($9.7\%$) on Cityscapes. Fig.˜4 shows linear probe segmentation result.
@@ 4.3.2 Depth Estimation
For depth estimation, we follow the approach in [featup] and train on pseudo-labels generated by the state-of-the-art Depth Anything V2 network [depth_anythingv2]. We report two standard metrics from the monocular depth estimation literature: root mean square error (RMSE) and $\delta_{1}<1.25$. The $\delta_{1}$ metric measures the percentage of pixels where the predicted depth $y$ is within 25% of the ground-truth $y^{*}$, formally defined as $\delta_{1}=max\left(\frac{y}{y^{*}},\frac{y^{*}}{y}\right)<1.25$. We train the linear probe for 5 epochs on the COCO training set, using a batch size of 4. Although JAFAR was not trained on this specific task, we observe that it reaches competitive scores, ranking second among the baselines. Notably, JAFAR outperforms both FeatUp and LiFT while also surpassing all task-dependent methods but ReSFU. Fig.˜4 shows linear probe depth estimation result.
@@ 4.3.3 Class Activation Maps Faithfulness
Following the approach in [featup], our method can be seamlessly integrated into explainability tools such as Class Activation Maps (CAMs). Despite recent advances, CAMs are still fundamentally limited by the low-resolution feature maps produced by standard vision encoders, which hinders their ability to localize fine-grained details. By upsampling the features, our method yields sharper and more informative explanations. To assess the quality of the resulting CAMs, we adopt standard evaluation metrics from the literature: Average Drop (A.D), Average Increase (A.I), Average Gain (A.G), Coherency (Coh.), and Complexity (Cplx.).
@@
A.D ($\downarrow$) A.I ($\uparrow$) A.G ($\uparrow$) Coh. ($\uparrow$) Cplx. ($\downarrow$) ADCC ($\uparrow$) Training-free      Bilinear 19.0 18.5 3.4 88.8 59.9 61.7 (-11.6)      Large Image (x8) 48.8 12.8 2.5 67.9 38.0 59.5 (-13.8)      Strided 19.8 15.1 3.5 85.9 54.2 65.3 (-8.0) Task-Dependent      CARAFE [carafe] 49.9 4.8 1.0 66.9 34.9 59.7 (-13.6)      SAPA [sapa] 8.5 32.7 4.1 96.5 69.6 55.4 (-17.9)      Dysample [dysample] 17.8 20.0 3.8 90.2 60.7 61.6 (-11.7)      ReSFU [resfu] 14.5 24.0 3.7 92.1 64.2 59.4 (-13.9) Task-Agnostic      FeatUp [featup] 15.3 24.0 4.3 91.6 58.2 64.3 (-9.0)      LiFT [lift] 66.9 8.7 2.3 65.2 9.4 53.0 (-20.3)      JAFAR 17.4 30.9 6.5 91.4 44.1 73.3
In particular, A.D, A.I, and A.G measure how sensitive the classifier’s output is to the most salient regions of the input—an effective CAM should highlight areas that, when masked, lead to a notable change in classification confidence. Since each of these metrics captures only a single aspect of CAM quality, we also report the ADCC score—an aggregate metric proposed in [adcc] that provides a more holistic evaluation. Additional details are provided in Sec.˜A.1. As illustrated qualitatively in Fig.˜4, JAFAR generates sharper and more semantically accurate CAMs compared to all baselines. While training-free methods don’t help to recover important regions, task-dependent approaches typically produce blurrier and less precise maps. Quantitative results in Tab.˜2 further support this, with JAFAR achieving the highest score on the aggregate ADCC metric—outperforming the second-best method by 8 points, a relative improvement of 12.5%.
@@ 4.3.4 Zero-Shot Open-Vocabulary Segmentation
We further evaluate our method on a zero-shot open-vocabulary segmentation task, following the setup from [maskclip], where class labels from the dataset serve as textual inputs and predictions are made by selecting the class with the highest similarity score (argmax). Using a CLIP-ViT-B/16 backbone, this approach is entirely training-free, as it does not require a learned probing head. Results show that JAFAR significantly outperforms all baselines, with particularly strong improvements on Pascal VOC. Despite the increased difficulty of ADE20K, which includes 150 classes, our method still achieves the highest performance in both mIoU and accuracy. We report only FeatUp among the task-agnostic baselines, as it is the second-best performing method.
Table 3: Zero-Shot Open-Vocabulary Evaluation. Using MaskCLIP [maskclip] for zero-shot open-vocabulary segmentation, JAFAR consistently improves performance, indicating strong alignment with the original features.
VOC ADE20K Cityscapes Upsampling mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) Nearest 24.13 30.80 9.33 24.65 19.66 50.27 Bilinear 27.87 35.27 11.03 27.78 21.56 53.21 Large Image ($\times 2$) 23.24 32.16 8.08 24.94 21.91 52.22 FeatUp [featup] 32.27 39.78 13.03 33.28 24.76 60.11 JAFAR 35.70 44.93 13.61 33.28 25.26 61.73
@@ 4.3.5 Bird’s-Eye View Segmentation
Upsampling SimpleBeV [harley2023simple] PointBeV [pointbev] BeVFormer [bevformer] Low-Res 31.75 34.89 33.72 Bilinear 33.67 36.01 34.18 FeatUp 33.95 35.38 34.01 JAFAR 36.59 37.20 36.54
Finally, we studied the impact of our upsampler in a complex training pipeline. The task, evaluated on nuScenes [nuscenes], takes several images taken from cameras as input and consists on outputting the bird’s-eye view (BeV) segmentation map. In our setup, we used a frozen DINOv2 [dinov2] backbone and trained the rest of the architecture, namely, the upsampler, the BeV encoder, and the segmentation head. This task is particularly challenging, as the model must learn to map features from the image plane to the BeV plane. To ensure a fair comparison, we also trained the architecture without an upsampler, using lower-resolution input images $(224\times 476)$. We adopted the optimization hyperparameters from PointBeV [pointbev], adjusting the batch size to 1 and training for 100 epochs. Our results show that using an upsampler consistently improves predictions, regardless of the architecture employed—SimpleBev [harley2023simple], PointBeV [pointbev], or BevFormer [bevformer]. Notably, performance improves significantly when using JAFAR as the upsampler, with mIoU gains up to +5+5 points.
@@ 4.4 Ablations
To evaluate the benefit of deriving both queries and keys from a shared image encoding, we compare in Tab.˜5 several strategies to obtain keys. In the Linear Projection baseline, keys are obtained by applying a linear layer to the vision encoder’s low-resolution features $F_{lr}$, to match JAFAR’s embedding dimension. In the Concatenation baseline, semantics is injected via a direct concatenation of $F_{lr}$ and preliminary keys $\tilde{K}$.
@@
Ablation Type / Setting Semantic Segmentation VOC ADE20K Cityscapes mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) mIoU ($\uparrow$) Acc ($\uparrow$) Keys Strategy    Linear Projection 80.02 (-4.42) 94.87 (-1.41) 37.87 (-2.62) 73.22 (-1.70) 52.45 (-9.02) 90.80 (-2.62)    Concatenation 83.13 (-1.27) 95.94 (-0.34) 40.06 (-0.43) 74.56 (-0.36) 58.70 (-2.77) 92.65 (-0.77)    w/o SFT 83.25 (-1.19) 95.93 (-0.35) 39.62 (-0.87) 74.32 (-0.60) 56.53 (-4.94) 92.19 (-1.23)    w/ SFT 84.44 96.28 40.49 74.92 61.47 93.42 Attention Heads    $n=1$ 84.13 (-0.31) 96.21 (-0.07) 40.15 (-0.34) 74.79 (-0.13) 60.94 (-0.53) 93.32 (-0.10)    $n=2$ 84.27 (-0.17) 96.27 (-0.01) 40.42 (-0.07) 74.95 (+0.03) 61.19 (-0.28) 93.42 (-0.00)    $n=4$ 84.44 96.28 40.49 74.92 61.47 93.42    $n=8$ 83.82 (-0.62) 96.13 (-0.15) 40.07 (-0.42) 74.20 (-0.72) 60.56 (-0.91) 93.33 (-0.09)
In comparison, the Linear projection baseline shows a significant performance drop, and SFT consistently outperforms the concatenation approach. Increasing the number of attention heads up to 4 further enhances performance by producing more robust upsampling kernels through averaged post-softmax scores. Beyond this point, however, the benefits reverse: the per-head dimensionality becomes too low to support effective alignment, while the computational cost increases, ultimately degrading output quality.